OpenStudy (anonymous):

Can anyone explain to me how to find the slope-intercept form of 2x-3y=0?

OpenStudy (victoriasushchik):

SLOPE = -2/3 Y intercept= -2 Always change the given equations into Y=M(sope)X+B(y Intercept) To graph this take the y=-2/3x-2, -2 would go on the Y line ( up and down part ) since it is the Y intercept and then take it down by -2 ( the -2 from the slop (-2/3) and take it to the right by 3. ( because slope is rise over run, and -2 is the rise ( up and down ) and 3 is the run and is always going to go positive ( to the right ) I love how that * guy a few posts below me copy and pasted my answer because he was too lazy to do it himself lol. And Eric W a standard form of a line is not y=mx+b a standard for is ax+by=c, y=mx+b is the slope intercept form of a line,

OpenStudy (anonymous):

Ok thank you very much for your help.

OpenStudy (victoriasushchik):

your welcome hon! remember medal me and fan!